Slope of KE versus position vector gives rate of change of ???

Grade:11

1 Answers

Vikas TU
7 years ago
K = 0.5mv^2
or
K = 0.5m*(dx/dt)^2
root(2K/m) = dx/dt
dx = root(2K/m)dt
integerate both sides,
x = root(2K/m)t
squaring both sides,
x^2 = (2K/m)t^2
K = (m/2)(x/t)^2